# A Lightweight 3D Anomaly Detection Method with Rotationally Invariant Features

### v1

@@ 4.5 Generalization of Our Framework RIF
Table 7: Results of RIF with other feature extractors on Anomaly-ShapeNet in terms of different evaluation metrics.
As shown in Table 7, the integration of RIF with other feature extractors greatly enhances anomaly detection performance. On Anomaly-ShapeNet, RIF with PCM achieved notable improvements over baseline (w/o PCM): PointNet (P-AUROC: +32.67%, O-AUROC: +17.33%), PointMLP (P-AUROC: +7.63%, O-AUROC: +17.82%), DGCNN (P-AUROC: +16.04%, O-AUROC: +19.70%), PointMAE (P-AUROC: +8.00%, O-AUROC: +9.83%), Stratified-Transformer (P-AUROC: +2.15%, O -AUROC: +5.43%), and PointTransformer-v3 (P-AUROC: +15.14%, O-AUROC: +11.83%).
For Real3D-AD, as shown in Table 8, RIF with PCM obtained significant performance improvement, particularly with PointNet (P-AUROC: +24.23%, O-AUROC: +8.80%), DGCNN (P-AUROC: +16.17%, O -AUROC: +3.15%), Stratified-Transformer (P-AUROC: +10.23%, O -AUROC: +3.14%), and Point Transformer-v3 (P-AUROC: +6.12%, O-AUROC: +2.30). These results demonstrate the strong adaptability of RIF to convolution-based (PointNet, DGCNN), MLP-based (PointMLP), and transformer-based (PointMAE, Stratified-Transformer, PointTransformer-v3) architectures, making it highly suitable for industrial 3D anomaly detection scenarios where available data exhibits a high degree of diversity and variability.
@@
Table 8: Results of RIF with other feature extractors in terms of different evaluation metrics.
4.6 Efficiency Comparison between Our Method and Existing Methods
@@
4.7 Potential Data Leakage Issues
ModelNet is used as a pre-training method in some baselines. To avoid potential data leakage, we utilized S3DIS [Armeni2016CVPR] for pre-training, and the results are shown in Table 9. Similar detection results from pre-training on a large indoor dataset validate that the proposed method does not suffer from data leakage.
@@ 4.7 Potential Data Leakage Issues
Table 9: Performance comparison on pre-training with different datasets.
5 Conclusion
@@
6 Acknowledgments
@@
7 Appendix
The Proof of the Point Coordinate Mapping (PCM) for Rotational Invariance.
7.1 Definitions
Definition A1: Given a point cloud $\bm{\mathcal{P}}$ = $\{\textbf{p}_{i}\}_{i=1}^{n}$ $\in$ $\mathbb{R}^{n\times 3}$ with each point $\textbf{p}_{i}$ = $(x_{i},y_{i},z_{i})^{\top}$, the centroid of $\bm{\mathcal{P}}$ is defined as:
$\textbf{c}=\frac{1}{n}\sum_{i=1}^{n}\textbf{p}_{i}.$ (10)
Definition A2: Given a point cloud $\bm{\mathcal{P}}$ = $\{\textbf{p}_{i}\}_{i=1}^{n}$ $\in$ $\mathbb{R}^{n\times 3}$, the relative coordinate vector of each point $\textbf{p}_{i}$($i\text{=}1,2,\ldots,n$) with respect to the centroid c is defined as:
$\textbf{r}_{i}=\textbf{p}_{i}-\textbf{c}.$ (11)
Definition A3: Let R $\in$ SO(3) be a matrix with the size of 3*3. R is a unitized rotation matrix if and only if:
1) $\textbf{R}\textbf{R}^{\top}=\textbf{I}_{3}$,
2) $Det(\textbf{R})=1$,
where SO(3) denotes the 3D rotation group, and $Det$ means the value of the determinant.
Definition A4: Given a rotation matrix R, the transformed relative coordinate of each point $\textbf{p}_{i}$ is defined as:
$\textbf{p}^{\prime}_{i}=\textbf{R}\textbf{r}_{i}+\textbf{c}.$ (12)
Definition A5: Given a point cloud $\bm{\mathcal{P}}$ = $\{\textbf{p}_{i}\}_{i=1}^{n}$ $\in$ $\mathbb{R}^{n\times 3}$, the three key vectors $\textbf{u}_{1},\textbf{u}_{2},\textbf{u}_{3}$ to define the relative coordinate system can be formalized as:
$\textbf{u}_{1}$ $=\textbf{p}_{\text{far1}}-\textbf{c},$ (13)
$\textbf{u}_{2}$ $=\textbf{p}_{\text{far2}}-\textbf{c},$
$\textbf{u}_{3}$ $=\textbf{p}_{\text{near}}-\textbf{c},$
where $\textbf{p}_{\text{far1}}$ is the farthest point from the centroid, $\textbf{p}_{\text{far2}}$ is the second farthest point from the centroid and is linearly independent to $\textbf{u}_{1}$, and $\textbf{p}_{\text{near}}$ is the nearest point to the centroid that makes the rank of the three selected vectors equal to 3, i.e., $Rank([\textbf{u}_{1};\textbf{u}_{2};\textbf{u}_{3}])=3$.
Definition A6: Given three key vectors $\textbf{u}_{1},\textbf{u}_{2},\textbf{u}_{3}$, the orthogonalized vectors after Schmidt orthogonalization can be defined as:
$\textbf{e}_{1}$ $=\frac{\textbf{u}_{1}}{\|\textbf{u}_{1}\|},$ (14)
$\textbf{e}_{2}$ $=\frac{\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{e}_{1}}{\|\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{e}_{1}\|},$
$\textbf{e}_{3}$ $=\frac{\textbf{u}_{3}-(\textbf{u}_{3}\cdot\textbf{e}_{1})\textbf{e}_{1}-(\textbf{u}_{3}\cdot\textbf{e}_{2})\textbf{e}_{2}}{\|\textbf{u}_{3}-(\textbf{u}_{3}\cdot\textbf{e}_{1})\textbf{e}_{1}-(\textbf{u}_{3}\cdot\textbf{e}_{2})\textbf{e}_{2}\|}.$
Definition A7: Given a point cloud $\bm{\mathcal{P}}$ = $\{\textbf{p}_{i}\}_{i=1}^{n}$ $\in$ $\mathbb{R}^{n\times 3}$, the Point Coordinate Mapping (PCM) with rotational invariance and the relative coordinate point after mapping are defined as:
$\bm{\textbf{M}}:\bm{\textbf{P}}\rightarrow\bm{\textbf{P}}^{*},$ (15)
$\textbf{p}_{i}^{*}=(\textbf{p}_{i}-\textbf{c})[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]^{\text{T}},$ (16)
where $\textbf{p}_{i}^{*}$ represents the point with rotational invariance after being processed by PCM.
7.2 Propositions
Proposition A1: The coordinate representation of the centroid of the point cloud $\bm{\mathcal{P}}$ is rotationally invariant.
Proof.
Given a rotation matrix R satisfied Definition A3, the centroid $\bm{\textbf{c}}^{\prime}$ of the rotated point cloud $\bm{\mathcal{P}^{\prime}}$ = $\{\textbf{p}^{\prime}_{i}\}_{i=1}^{n}$ satisfied Definition A4 has the following property:
$\textbf{c}^{\prime}$ $=\frac{1}{n}\sum_{i=1}^{n}\textbf{p}_{i}^{\prime}$ (17)
$=\frac{1}{n}\sum_{i=1}^{n}\left(\textbf{R}(\textbf{p}_{i}-\textbf{c})+\textbf{c}\right)$
$=\textbf{R}\left(\frac{1}{n}\sum_{i=1}^{n}(\textbf{p}_{i}-\textbf{c})\right)+\textbf{c}$
$=\textbf{c}.$
Thus, the centroid of the point cloud is rotationally invariant. The proof of Proposition 1 is complete. ∎
Proposition A2: the orthogonalized vectors $\textbf{e}_{1}^{\prime},\textbf{e}_{2}^{\prime},\textbf{e}_{3}^{\prime}$ are simply the original orthogonalized vectors $\textbf{e}_{1},\textbf{e}_{2},\textbf{e}_{3}$ transformed by the rotation matrix R.
Proof.
Given the relative position vectors $\textbf{r}_{i}(i=1,2,\ldots,n)$ satisfied Definition A2 and R $\in$ $\mathrm{SO(3)}$ satisfied Definition A3, it has:
$\textbf{r}^{\prime}_{i}$ $=\textbf{p}_{i}^{\prime}-\textbf{c}^{\prime}$ (18)
$=\textbf{R}(\textbf{p}_{i}-\textbf{c})$
$=\textbf{R}\textbf{r}_{i}.$
Given three key vectors $\textbf{u}_{1},\textbf{u}_{2}$, and $\textbf{u}_{3}$ satisfied Definition A5 and R $\in$ $\mathrm{SO(3)}$ satisfied Definition A3, these vectors after rotation can be represented as:
$\textbf{u}_{1}^{\prime}=\textbf{R}\textbf{u}_{1},$ (19)
$\textbf{u}_{2}^{\prime}=\textbf{R}\textbf{u}_{2},$
$\textbf{u}_{3}^{\prime}=\textbf{R}\textbf{u}_{3}.$
After Schmidt Orthogonalization, these vectors can be represented as:
$\textbf{e}_{1}^{\prime}$ $=\frac{\textbf{u}_{1}^{\prime}}{\|\textbf{u}_{1}^{\prime}\|}$ (20)
$=\frac{\textbf{R}\textbf{u}_{1}}{\|\textbf{R}\textbf{u}_{1}\|}$
$=\textbf{R}\left(\frac{\textbf{u}_{1}}{\|\textbf{u}_{1}\|}\right)=\textbf{R}\textbf{e}_{1},$
$\textbf{e}_{2}^{\prime}$ $=\frac{\textbf{u}_{2}^{\prime}-(\textbf{u}_{2}^{\prime}\cdot\textbf{e}_{1}^{\prime})\textbf{e}_{1}^{\prime}}{\|\textbf{u}_{2}^{\prime}-(\textbf{u}_{2}^{\prime}\cdot\textbf{e}_{1}^{\prime})\textbf{e}_{1}^{\prime}\|}$ (21)
$=\frac{\textbf{R}\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{R}\textbf{e}_{1}}{\|\textbf{R}(\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{e}_{1})\|}$
$=\textbf{R}\left(\frac{\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{e}_{1}}{\|\textbf{u}_{2}-(\textbf{u}_{2}\cdot\textbf{e}_{1})\textbf{e}_{1}\|}\right)=\textbf{R}\textbf{e}_{2}.$
Similarly,
$\textbf{e}_{3}^{\prime}=\textbf{R}\textbf{e}_{3}.$ (22)
Thus, the orthogonalized vectors $\textbf{e}_{1}^{\prime},\textbf{e}_{2}^{\prime},\textbf{e}_{3}^{\prime}$ are simply the original orthogonalized vectors $\textbf{e}_{1},\textbf{e}_{2},\textbf{e}_{3}$ transformed by the rotation matrix R. The proof of Proposition 2 is complete.
Proposition A3: Give a rotation matrix R $\in$ $\mathrm{SO(3)}$ and point cloud $\bm{\mathcal{P}}$ = $\{\textbf{p}_{i}\}_{i=1}^{n}$ $\in$ $\mathbb{R}^{n\times 3}$ , the pre-rotation and post-rotation representations of each point $\textbf{p}_{i}\text{=}(x_{i},y_{i},z_{i})^{T}(i\text{=}1,2,\ldots,n)$ being processed by PCM satisfies:
$(x_{i},y_{i},z_{i})_{\text{pre-rotation}}\equiv(x_{i},y_{i},z_{i})_{\text{post-rotation}}.$ (23)
Proof.
Given the original point cloud processed by PCM satisfied Definition A7, i.e., $\textbf{p}_{i}^{*}=(\textbf{p}_{i}-\textbf{c})\cdot[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]^{\top}$. For the rotated point cloud $\textbf{p}_{i}^{\prime}=\textbf{R}(\textbf{p}_{i}-\textbf{c})+\textbf{c}$ satisfied Definition A4, the computation of its PCM coordinates can be formed as follows:
1) Obtain the relative coordinate centroid (Proposition A1): $\textbf{c}^{\prime}=\textbf{c}$.
2) Generate the orthogonal basis vectors (Proposition A2): The rotated basis vectors satisfy $\textbf{e}_{1}^{\prime};\textbf{e}_{2}^{\prime};\textbf{e}_{3}^{\prime}]=\textbf{R}[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]$.
3) Perform PCM mapping for rotated points:
$\textbf{p}_{i}^{*^{\prime}}$ $=\textbf{R}(\textbf{p}_{i}-\textbf{c})\cdot\left(\textbf{R}[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]\right)^{\top}$ (24)
$=\textbf{R}\textbf{R}^{T}(\textbf{p}_{i}-\textbf{c})\cdot[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]^{\top}$
$=(\textbf{p}_{i}-\textbf{c})\cdot[\textbf{e}_{1};\textbf{e}_{2};\textbf{e}_{3}]^{\top}$
$=\textbf{p}_{i}^{*}.$
Thus, PCM is rotationally invariant. The proof of Proposition 3 is complete. ∎
In all, the PCM-processed coordinates $\textbf{p}_{i}^{*}$ are invariant under any rotation $\textbf{R}\in\mathrm{SO(3)}$, as the transformed coordinates $\textbf{p}_{i}^{*^{\prime}}$ equal $\textbf{p}_{i}^{*}$. Thus, PCM has the property of rotational invariance.

### v2

@@ IV-E Generalization of Our Framework RIF
TABLE VII: Results of RIF with other feature extractors on Anomaly-ShapeNet in terms of different evaluation metrics.
As shown in Table VII, the integration of RIF with other feature extractors greatly enhances anomaly detection performance. On Anomaly-ShapeNet, RIF with PCM achieved notable improvements over baseline (w/o PCM): PointNet (P-AUROC: +32.67%, O-AUROC: +17.33%), PointMLP (P-AUROC: +7.63%, O-AUROC: +17.82%), DGCNN (P-AUROC: +16.04%, O-AUROC: +19.70%), PointMAE (P-AUROC: +8.00%, O-AUROC: +9.83%), Stratified-Transformer (P-AUROC: +2.15%, O -AUROC: +5.43%), and PointTransformer-v3 (P-AUROC: +15.14%, O-AUROC: +11.83%).
For Real3D-AD, as shown in Table VIII, RIF with PCM obtained significant performance improvement, particularly with PointNet (P-AUROC: +24.23%, O-AUROC: +8.80%), DGCNN (P-AUROC: +16.17%, O -AUROC: +3.15%), Stratified-Transformer (P-AUROC: +10.23%, O -AUROC: +3.14%), and Point Transformer-v3 (P-AUROC: +6.12%, O-AUROC: +2.30). These results demonstrate the strong adaptability of RIF to convolution-based (PointNet, DGCNN), MLP-based (PointMLP), and transformer-based (PointMAE, Stratified-Transformer, PointTransformer-v3) architectures, making it highly suitable for industrial 3D anomaly detection scenarios where available data exhibits a high degree of diversity and variability.
@@
TABLE VIII: Results of RIF with other feature extractors in terms of different evaluation metrics.
IV-F Efficiency Comparison between Our Method and Existing Methods
@@
IV-G Potential Data Leakage Issues
ModelNet is used as a pre-training method in some baselines. To avoid potential data leakage, we utilized S3DIS [Armeni2016CVPR] for pre-training, and the results are shown in Table IX. Similar detection results from pre-training on a large indoor dataset validate that the proposed method does not suffer from data leakage.
@@ IV-G Potential Data Leakage Issues
TABLE IX: Performance comparison on pre-training with different datasets.
V Conclusion
@@
Limitation: Although the proposed approach has yielded promising results, its patch-based representation imposes inherent limitations on achieving fine-grained, point-level predictive accuracy. As part of our future work, we aim to extend rotation-invariant reconstruction paradigms to enable more precise, point-wise anomaly localization.
VI Acknowledgments